Your car travels at 75 mph on the interstate. You drive at this rate for 90 minutes. If your car tires have a diameter of 20 inches, how many times will each tire revolve while driving those 90 minutes?
Answer by lwsshak3(11628)

Your car travels at 75 mph on the interstate. You drive at this rate for 90 minutes. If your car tires have a diameter of 20 inches, how many times will each tire revolve while driving those 90 minutes?
**
Distance car traveled=speed*travel time=75 mi/hr*90 min=
75*1.5=112.5 mi*5280 ft/mi*12 in/ft=7.128*10^6 in
circumference of tire=π*diameter=π*20 in
Times will each tire revolve while driving those 90 minutes
=distance traveled/circumference of tire
=7.128*10^6/π*20=(3.564/π)*10^5